OpenStudy (anonymous):

What is a set of real or complex number or elements arranged in rows or columns to form a rectangular array. (a)set (b)matrix (c)determinant (d)real number

hartnn (hartnn):

thats would be the definition of matrices just 'set' is a very general term determinant is not a set, it has a value.

OpenStudy (unklerhaukus):

the last word in the question is a synonym for the answer

OpenStudy (anonymous):

U haven't answered my question

hartnn (hartnn):

oh, i forgot to mention that matries is plural for matrix

hartnn (hartnn):

**matrices

OpenStudy (unklerhaukus):

array=matrix

OpenStudy (anonymous):

Matrix is the correct ans

hartnn (hartnn):

you got it:)

OpenStudy (anonymous):

Solve the set linear equations by guassian elimination method :a+2b+3c=5,3a-b+2c=8,4a-6b-4c=-2.find c (a) 4 (b) 5 (c) 9 (d) 10

OpenStudy (debbieg):

Welcome to Open Study @ginika ! Please post each problem as a separate question. You'll have a better chance of being helped with it. Also, you should review the site Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

hump day woot woot.